ARGOS Engineering half night 20180419

02:15 Open and ready to send preset
  • Will start with LAT test.
  • In parallel I will try to collect TT LUT Data
  • Presets
    • Steve set pointing
    • Doug send ACTIVE Preset to on-axis star: collimate
    • Doug send ACE-TRACK Preset to close the loop

02:20 MOS error on LUCI2 when requesting to remove mask from focal plane
  • Called Dave T.

02:28 First star will be #17, AO975

02:38 Binoc Preset to AO975

03:19 Closed loop on SX
  • Can not get AO Supervisor to collect LUCI images on either side
  • Dark 01
  • Align mask on SX

03:37 SX
  • First image LUCI1 4, TN 033737 STAGES were powered on

03:39 DX
  • Dark 01
  • First image

03:45 SX
  • First image LUCI1 9, TN 034610

03:59 DX
  • First image LUCI2 10, TN 040122

-- %USERSIG{DougMiller - 2018-04-19}%
Topic revision: r2 - 19 Apr 2018, DougMiller
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback